Dawn's mother is 4 years older than twice dawn's age. The difference of their age is 22 years. Find their ages
Answer by macston(5194)

M=mother's age; D=Dawn's age
M=2D+4yrs
M-D=22yrs Substitute for M
2D+4yrs-D=22yrs Subtract 4yrs from each side.
D=18yrs ANSWER 1: Dawn is 18 years old.
M=2D+4yrs=2(18yrs)+4yrs=36yrs+4yrs=40yrs ANSWER 2: Dawn's mother is 40 years old.
CHECK:
M-D=22yrs
40yrs-18yrs=22yrs
22yrs=22yrs